OpenStudy (anonymous):

Algebra 1 World problems systems of equations. Manny, Moe, and jack drove 168 miles. If Manny drove twice as far as Jack and Moe Drove three times as far as Jack, how far did Jack drive? Can you explain how to do this so I can do this for the other 3 problems please help And I'd appreciate if you'd help. T^.^

OpenStudy (anonymous):

Really need help

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

hang on let me see if something works real quick

hero (hero):

x + y + z = 168 x = 2z y = 3z 2z + 3z + z = 168

OpenStudy (anonymous):

If it was 168 why did my teacher put 60 instead. o.o

hero (hero):

It says they drove a total of 168 miles

OpenStudy (anonymous):

Should I believe you guys more ? Or my teacher is incorrect

hero (hero):

That's the first sentence

OpenStudy (anonymous):

Okay the answers ends up being -- Jack drove 28 miles. How I did it: A (manny) + B (moe) + C (jack) = 168. The problem states that Manny drove twice as far as Jack so A = 2C. Moe drove three times as far as Moe, so B = 3C. You now fill A and B in creating this equation: 2C + 3C + C = 168. Combine like terms making it 6c = 168, divide by 6, there's your answer.

hero (hero):

@mariannu, you are confused

OpenStudy (anonymous):

I know But this is what he did x+x+1+x+2=60

hero (hero):

Well, that's wrong

OpenStudy (anonymous):

he got an answer of x=9

hero (hero):

Are you sure that's a teacher or an imposter?

OpenStudy (anonymous):

Maybe he did not realize he put 60 because sometimes he would do the equation to fast and he would make an error sometimes. NOt often

hero (hero):

No, the entire setup is wrong. I think you may have confused which problem the teacher was solving.

OpenStudy (anonymous):

I'm still confused.

hero (hero):

Well, that's probably your teacher's fault

OpenStudy (anonymous):

hmm ok XD I will just do it your way I guess. XD Thanks though.

hero (hero):

What's your teacher's phone number? I will call him and explain to him what the correct answer is.

OpenStudy (anonymous):

I wonder where the hell he got that equation. o.o Usually can keep up with his fast paste.XD

OpenStudy (anonymous):

I will just go to him tomorrow morning and tell him about it And thank you. ^.^ Both of you guys.

OpenStudy (anonymous):

No its ok. No need too. I can tell him, but thanks though for your generosity . ^.^

OpenStudy (anonymous):

And plus I don't know his number anyways.

hero (hero):

Maybe your teacher copied the wrong solution from the answer book.

OpenStudy (anonymous):

hmm probably. but thanks ^.^